Pairs of markings a set distance apart are made on highways so that police can detect drivers exceeding the speed limit. Over a fixed​ distance, the speed R varies inversely with the time T. In one particular pair of​ markings, ris 71 mph when t is 2 seconds. Find the speed of a car that travels the given distance in 3 seconds.


1
Expert's answer
2022-04-08T10:04:41-0400

The distance is

"D=RT."

Since distance is fixed:

"R_1T_1=R_2T_2, \newline
R_2=R_1\frac{T_1}{T_2}=71\cdot\frac{2}{3}= 47.3\space mph."